OpenStudy (anonymous):

Randy and Sandy are 500 km apart. Randy starts driving towards Sandy at 1:00 and maintains a speed of 90 km/h. Sandy starts driving towards Randy at 2:00 and maintains a speed of 105km/h. How far has Randy traveled when they meet?

OpenStudy (anonymous):

I think i need to find the time each has traveled....?

OpenStudy (anonymous):

to then determine the distance...since i already have the rate at which each was traveling

OpenStudy (anonymous):

just not sure how to do that

OpenStudy (anonymous):

sandy's time is 1 less than randys time

OpenStudy (amistre64):

randy is an hour ahead and has cut off 90 km of his distance

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so sandy can be represented as t-1

OpenStudy (amistre64):

so start them heading towards each other at a distanceof 500-90

OpenStudy (anonymous):

ok they are heading towards each other at a distance of 410

OpenStudy (amistre64):

yay!! how much time do they travel combined in 1 hour?

OpenStudy (anonymous):

195

OpenStudy (amistre64):

195 km right...good now how many hours is that till they are almost together?

OpenStudy (amistre64):

410/195

OpenStudy (anonymous):

2

OpenStudy (amistre64):

after 2 hours how much distance is left?

OpenStudy (anonymous):

2.1 to be exact

OpenStudy (anonymous):

20 km

OpenStudy (amistre64):

at a combined spped of 195km per hour, how long does it take them to cover 20 km?

OpenStudy (anonymous):

195/20

OpenStudy (anonymous):

other way around?

OpenStudy (anonymous):

.10

OpenStudy (amistre64):

and 1/10 of an hour is..... 6 minutes right?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

lets make sure this last bit is right....let me go over it...

OpenStudy (anonymous):

wait, if it takes them both 2.1 hours to travel the distance

OpenStudy (amistre64):

so add an hour to that and use it to calclate randys distance

OpenStudy (anonymous):

nevermind i think thats wrong

OpenStudy (amistre64):

they meet up at the same point :) when they meet, the distance is traveled

OpenStudy (anonymous):

if 410 km is traveled combined in 2.1 hours

OpenStudy (anonymous):

do I then make an equation of 2.1= t + (t-1)

OpenStudy (amistre64):

yes, when they started together they had a distance of 410 km to cover together, they meet after "so much time" passes".

OpenStudy (amistre64):

90t =-105t right? maybe not...but it looks good

OpenStudy (anonymous):

im confused now

OpenStudy (amistre64):

195t - 410 = 0 is what we want I think

OpenStudy (anonymous):

i feel like im making this more complicated than it is

OpenStudy (amistre64):

at 195 km/hour and at "t" hours they cover 410 miles, so the distance is 0 195t-410 = 0

OpenStudy (amistre64):

t = 410/195

OpenStudy (anonymous):

but thats t for both of them

OpenStudy (amistre64):

Randy has traveled for an extra hour so 90(t+1) = distance traveled by randy

OpenStudy (anonymous):

wow i feel stupid

OpenStudy (anonymous):

haha

OpenStudy (amistre64):

lol :)

OpenStudy (anonymous):

you really are superhero lol

OpenStudy (amistre64):

279.23 km ;) I wanted was hoping for a cool title like "turkey club on rye" but nooooooo.....

OpenStudy (anonymous):

i got the answer. HAHAHA